# Can We Verify Step by Step for Incorrect Answer Detection?

### v2

@@ 3.3 Experimental Setup
We will detail experimental setup in this section.
@@
Baseline. From Wang et al. (2022); Zhao et al. (2023a), ADS can act as the baseline in our setting.
Metrics. We will employ the F1 score and AUC-PR as our evaluation metrics, with precision and recall also included as supplementary measures. As we aim to detect potential incorrect answers, False labels are treated as positive. Therefore, true positives are examples with $L=F$ and $\hat{L}=F$.
PDS Implementation. For $\text{PPSS}(\cdot,\cdot)$, we adopt the SAUMMAC zero-shot model (Laban et al., 2022) that trained on MNLI (Williams et al., 2018), which is a consistency detection method for text summarization that leverages the out-of-the-box natural language inference model to compute pairwise entailment score. As suggested by Laban et al. (2022), probabilities of sentence-level entailment and contradiction are used. To detect all potential inconsistencies in any sentence, we use $min$ operation to aggregate across sentence pairs:
$\text{PPSS}(A,B)$ $=\min_{j}\max_{i}\bigl{(}ent(a_{i},b_{j})$ (2)
$\quad-con(a_{i},b_{j})\bigr{)},$
where $a_{i},b_{j}$ are sentences of $A,B$ correspondingly, $ent$ represents the probability of entailment, $con$ is the probability of contradiction, the $min$ is taken over all sentences of $B$, and $max$ is taken over all sentences of $A$. Note that the PPSS metric lacks symmetry; however, this is not a problem in our implementation because the PSS calculation averages across all possible ordered response pairs.
Threshold Selection. As $\text{ADS}\in\{0,1,2,...,5\}$, threshold $H$ is set to be the midpoint $H=2.5$. Another reason for this selection is that $\text{ADS}>2.5$ means the majority agrees, which aligns the choice in Zhao et al. (2023a). The range of this PPSS spans in $(-1,1)$. A negative value implies the potential presence of conflicting information. Consequently, it is reasonable to set the threshold $H$ for the PDS to be 0. Our choice of threshold $H$ is not task-specific. We regard this as a benefit because held-out data for threshold tuning may not always be available in the real scenario. A detailed discussion is given in Appendix C.3.
@@
Table 1: PDS consistently outperforms ADS across all subsets in our R2PE benchmark in F1 scores (in %). The abbreviations GPT-3, GPT-instruct, GPT-3.5, Gemini, Mixtral, mistral correspond respectively to the LLMs text-davinci-003, GPT-3.5-turbo-instruct, GPT-3.5-turbo, Gemini Pro, Mixtral-8x7b, and mistral-medium. Results on FEVER, HotpotQA, and 2WikiMultihop from mistral-medium are not reported, as discussed in Section 2. The last row and column are the average improvement of PDS over ADS across datasets and LLMs respectively.
@@
GSM8K ADS 70.79 55.83 54.58 52.59 63.18 48.15 -
PDS 74.14 61.02 55.37 56.23 68.37 51.86 +3.66
MATH ADS 92.82 85.50 78.28 87.81 90.13 88.05 -
PDS 93.83 86.77 80.84 89.49 91.82 89.66 +1.64
StrategyQA ADS 49.56 51.42 49.36 49.85 59.55 56.27 -
PDS 49.97 54.49 52.87 51.18 62.08 67.22 +3.63
Play ADS 44.37 48.93 49.90 46.00 72.68 50.44 -
PDS 46.14 48.01 51.31 47.66 72.84 51.86 +0.92
Physics ADS 39.94 40.60 48.63 79.10 53.05 52.58 -
PDS 46.85 48.58 61.31 78.50 56.99 60.66 +6.50
FEVER ADS 57.02 49.72 48.57 54.02 67.31 - -
PDS 62.19 51.41 49.00 55.43 66.91 - +1.66
HotpotQA ADS 89.71 79.14 83.99 90.15 74.89 - -
PDS 92.61 83.21 86.21 91.47 79.41 - +3.01
2WikiMultihop ADS 83.26 75.52 48.12 79.37 59.51 - -
PDS 86.72 79.80 54.37 80.20 58.94 - +2.38
avg - +3.12 +3.33 +3.74 +1.41 +1.91 +5.15 +2.97
@@
Table 3: Ablation study of PDS. All metrics are F1 scores (in %). Top: subsets related to HotpotQA on R2PE. Bottom: subsets related to 2WikiMultihop on R2PE.
@@ 4 Results and Analysis
4.1 PDS Substantially Outperforms ADS
As shown in Table 1, across all 45 subsets in our R2PE, PDS consistently improves ADS performance, yielding an average improvement of $5.10\%$ in terms of F1. Although ADS demonstrates greater precision, it suffers from significantly lower recall, indicating its propensity to overlook numerous examples with $L=F$. In contrast, PDS adeptly balances precision and recall, culminating in superior overall performance relative to ADS (Table 15 and 16). From Table 2, PDS improves ADS in terms of AUC-PR in the majority of cases and yields comparable results (less than $1\%$ decrease) in the remaining ones. On average, PDS results in $3.01\%$ absolute improvement in AUC-PR. Notably, ADS is a fairly good baseline on GSM8K, MATH, HotpotQA, and, 2WikiMultihop to verify the correctness of answers given by CoT self-consistency, which is a widely adopted approach to improve reasoning abilities of LLMs. Our findings demonstrate that PDS can further augment the performance of ADS. For discriminative tasks (StrategyQA, Play, Physics, and Play), ADS has relatively low F1 scores and AUC-PR. This may be attributed to the fixed set of answers (e.g., yes or no) in these tasks, which inherently exhibit higher consistency, making verification based solely on answer agreement challenging. In contrast, PDS can relieve this issue by evaluating reasoning processes.
@@ 4.1 PDS Substantially Outperforms ADS
To meticulously evaluate the efficacy of the PDS, we delve into the (HotpotQA from GPT-4) subset, partitioning it into two sets: one where answers fail to reach a consensus greater than half, and its counterpart. We then label the examples with an ADS less than the threshold $H$ as $\hat{L}=F$, and all the others as $\hat{L}=T$. As shown in Figure 3 (left), while the ADS proficiently segregates the examples with $L=T$, its discernment for the "False" category is less acute. On the contrary, the PDS (Figure 3 middle) strikes a better balance between precision and recall. Eliminating the implicit answer verification component in Equation (1) results in diminished precision (Figure 3 right, discussed in Section 4.2).
@@ 4.2 Ablation Study
To explore alternatives to PDS and understand the contribution of its components to the overall performance, we have carried out an extensive ablation study involving the following variants:
@@
PDS w/o ans. The original PSS considers all $PPSS(R_{i},R_{j})$, where each response $R_{i}$ equals a reasoning path $r_{i}$ and an answer $a_{i}$. As the information from the answers $a_{i}$’ is already used to obtain the final answer $a$, we can discard them and only compute the entailment score among rationale pairs $PPSS(r_{i},r_{j})$, that is, $PDS_{w/o\>ans}=(\overline{ADS}+\sum_{i\neq j}PPSS(r_{i},r_{j}))/2$.
@@
PDS-HaloCheck;PDS-SelfCheckNLI. HaloCheck (Elaraby et al., 2023) and SelfCheckNLI (Manakul et al., 2023) can be alternatives to PSS, denoted by PDS-halocheck and PDS-selfcheckNLI, respectively. Discussion of the difference between these two against our PSS implementation is provided in Appendix C.4.
As shown in Table 3, our PDS implementation gains the best results.
@@ 4.3 PDS for Improving Downstream Performance
VE (Zhao et al., 2023a) post-edits reasoning process using external knowledge (Appendix D). VE first checks the answer agreement and then edits the reasoning chains based on a retriever if $ADS<H$. That is to say, VE selects potential incorrect samples by ADS. Our PDS can be incorporated into the VE framework, dubbed "VE + PDS".
Experiments are carried out on (HotpotQA from GPT-4) and (2WikiMultihop from GPT-4), utilizing ground-truth supporting contexts along with distractor paragraphs as the retrieved documents to negate the influence of the retrival system. Our baseline is VE, and the standard prompting method (Brown et al., 2020) and CoT-SC (Wang et al., 2022) are also included as a reference.
Figure 4 illustrates that integrating our method with the VE framework leads to enhanced accuracy for both datasets. Specifically, VE+PDS leads to a notable $2.9\%$ absolute increase in accuracy for HotpotQA, and $2.8\%$ for 2WikiMultihop. Given that the retrieval system in our experiments is assumed to operate with near-perfection, the observed improvements in performance are largely attributed to the more effective selection of samples, achievable through the substitution of ADS with PDS. This success is likely a consequence of the PDS capability to supervise processes across multiple reasoning chains, leading to better selection of all potential incorrect predictions.
@@ 5 Related Work
Extensions of CoT. CoT (Wei et al., 2022) encourages LLMs to detail their reasoning steps with few-shot examples. Wang et al. (2022) advance CoT by substituting greedy decoding with sampling techniques. Moreover, various strategies (Chen et al., 2022; Gao et al., 2023; Zhang et al., 2023a) prompt LLMs to express rationales as Python codes. The search for high-quality exemplars has led to both automated (Kojima et al., 2022; Zhang et al., 2022; Fu et al., 2022; Wan et al., 2023; Xu et al., 2023; Wang et al., 2023; Zhao et al., 2023b; Crispino et al., 2023) and hybrid (Shao et al., 2023; Shum et al., 2023; Diao et al., 2023; Zou et al., 2023) demonstration selection. To enhance LLM reasoning across various tasks, novel prompt adaptations (Yao et al., 2022, 2023; Besta et al., 2023; Ding et al., 2023; Zheng et al., 2023; Zhang et al., 2023b) have been developed. For complex mathematical problems and knowledge-intensive tasks, PHP (Zheng et al., 2023), cumulative reasoning (Zhang et al., 2023b), and verify-and-edit (VE) (Zhao et al., 2023a) have been crafted. Li et al. (2023) further extend VE by incorporating diverse knowledge sources. While most research prioritizes task performance metrics, our study poineers the exploration of the connection between reasoning processes and outcomes, aiming to verify final predictions by analyzing reasoning chains.
Rationale Quality Evaluation. Rationale analysis is key for understanding AI performance and limitations(Golovneva et al., 2022; Prasad et al., 2023; He et al., 2023). Rationale evaluation methods are either reference-based, comparing against a gold standard (Clinciu et al., 2021; Welleck et al., 2022; Saparov and He, 2022), or reference-free, using metrics like ROSCOE (Golovneva et al., 2022). While they link high-quality reasoning to better task performance, the direct impact of reasoning quality on prediction accuracy is under-researched. Our study fills this gap with a detailed quantitative analysis of this relationship.
Uncertainty Estimation. A range of methods measure LLM uncertainty by token probabilities (Kuhn et al., 2023; Malinin and Gales, 2020; Kadavath et al., 2022; Fomicheva et al., 2020); however, they are inapplicable to R2PE, which does not furnish probabilities for responses. Moreover, they concentrate on short answer sequences, aligning more closely with the standard prompting (Brown et al., 2020). In contrast, our work verifies the answer veracity of CoT, which encompasses both the reasoning chains and answer sequences. Different from Kuhn et al. (2023), the entailment model in our work is used to measure the informational similarities between two responses, where the problem setting and detailed implementation are distinct.
@@ 6 Conclusion
To shed more light on the connection between the quality of the reasoning steps and the final outcome, we present R2PE, the first benchmark that quantitatively analyzes whether we can validate the answer veracity by evaluating the reasoning chains on a variety of reasoning tasks across five different domains. Six different LLMs are used in the creation of R2PE to guarantee generalizability of our findings. We develop the PDS framework by ingegrating multiple rationales’ information to predict the falsehood of the final predictions, which significantly boosts the performance of answer checking on our R2PE benchmark. Furthermore, our approach can be easily combined with verify-and-edit framework to improve end-task performance.
Limitations
There are several limitations of our work. First, R2PE focuses exclusively on tasks with answers derived from a predefined set of options. Consequently, it does not cover open-ended generative tasks such as text summarization. Second, the creation of R2PE is based on the original CoT-SC implementation (Wang et al., 2022), which now has a plethora of extensions. Third, our PDS implementation only evaluates the relation among different reasoning chains, neglecting the internal logical consistency of individual chains. To address these limitations, future endeavors will involve the collection of additional examples that cover a broader spectrum of task types into R2PE, including those that require long-text generation. Moreover, we intend to use various extensions of self-consistency methods to enrich the R2PE benchmark. In addition, a more holistic assessment of the quality of reasoning chains is left for future work, and it is possible to integrate our method with other more advanced rationale editing techniques.
Ethics Statement
Our R2PE is designed as a benchmark for assessing the falsehood of final outputs produced by LLMs across various domains and different LLMs. We will make R2PE available under MIT license.
Our PDS offers a novel approach to evaluating the answer veracity of LLMs by examining the consistency among multiple reasoning chains. The PDS is computationally efficient and can be integrated with additional error-correction methodologies to enhance task-specific performance. On the one hand, the ability to flag inaccurate LLM predictions will prove invaluable to users, offering insights into the reliability of LLM-generated content. Additionally, once an answer is identified as incorrect, additional methods can be used for correction or further scrutiny, thus mitigating the risk of hallucination or misinformation.
Our work represents a foundational step towards cultivating more advanced methods that both high precision and high recall. Our effort not only enhances and reliability of LLMs, but also significantly contributes to the broader discourse on ethical AI development and deployment.

### v4

@@ 3.3 Experimental Setup
We will detail the experimental setup in this section.
@@
Baselines. From Wang et al. (2022); Zhao et al. (2023), ADS can act as the baseline in our setting. We also include SelfCheckGPT (Manakul et al., 2023) and HaloCheck (Elaraby et al., 2023) as our baselines.
Metrics. We will employ the F1 score and AUC-PR as our evaluation metrics. As we aim to detect potential incorrect answers, False labels are treated as positive. Therefore, true positives are examples with $L=F$ and $\hat{L}=F$. We do not compare with SelfCheckGPT and HaloCheck in terms of F1 score, as they do not provide thresholds for prediction.
PDS Implementation. For $\text{PPSS}(\cdot,\cdot)$, we adopt the SAUMMAC zero-shot model (Laban et al., 2022),which is a consistency detection method for text summarization that leverages the out-of-the-box natural language inference model to compute pairwise entailment score. As suggested by Laban et al. (2022), probabilities of sentence-level entailment and contradiction are used. To detect all potential inconsistencies in any sentence, we use $min$ operation to aggregate across sentence pairs:
$\text{PPSS}(R_{k},R_{l})$ $=\min_{j}\max_{i}\bigl{(}ent(a_{i},b_{j})\quad-con(a_{i},b_{j})\bigr{)},$ (2)
where $a_{i},b_{j}$ are sentences of two responses $R_{k},R_{j}$ correspondingly, $ent$ represents the probability of entailment, $con$ is the probability of contradiction, the $min$ is taken over all sentences of $R_{j}$, and $max$ is taken over all sentences of $R_{k}$. Note that the PPSS metric lacks symmetry; however, this is not a problem in our implementation because the PSS calculation averages across all possible ordered response pairs.
Threshold Selection. As $\text{ADS}\in\{0,1,2,...,5\}$, threshold $H$ is set to be the midpoint $H=2.5$. Another reason for this selection is that $\text{ADS}>2.5$ means the majority agrees, which aligns the choice in Zhao et al. (2023). The range of this PPSS spans in $(-1,1)$. A negative value implies the potential presence of conflicting information. Consequently, it is reasonable to set the threshold $H$ for the PDS to be 0. Following Li et al. (2023), the threshold of ADS is raised to 4.5 for discriminative tasks (yes/no, multiple-choice questions), and the PDS threshold is adjusted to 0.4 according to Equation (1). Our choice of threshold $H$ is not task-specific. We regard this as a benefit because held-out data for threshold tuning may not always be available in the real scenario. A detailed discussion is given in Appendix C.3.
@@
Table 1: PDS consistently outperforms ADS across all subsets in our R2PE benchmark in F1 scores (in %). The abbreviations GPT-3, GPT-instruct, GPT-3.5, Gemini, Mixtral, mistral correspond respectively to the LLMs text-davinci-003, GPT-3.5-turbo-instruct, GPT-3.5-turbo, Gemini Pro, Mixtral-8x7b, and mistral-medium. Results on FEVER, HotpotQA, and 2WikiMultihop from mistral-medium are not reported, as discussed in Section 2. The last row and column are the average improvement of PDS over ADS across datasets and LLMs respectively. The results of HotpotQA and 2WikiMultihop from GPT-3.5 are replaced with those from GPT-4 (see Appendix A.2).
@@
GSM8K SelfCheckGPT 53.87 35.67 30.35 27.07 40.30 29.29 36.09
HaloCheck 47.89 34.52 28.45 21.65 39.61 25.03 32.86
ADS 70.79 55.83 54.58 52.59 63.18 48.15 57.52
PDS 74.14 61.02 55.37 56.23 68.37 51.86 61.17
MATH SelfCheckGPT 89.04 74.55 55.60 73.46 85.51 82.96 76.85
HaloCheck 88.11 76.46 58.26 80.08 82.57 81.77 77.87
ADS 92.82 85.50 78.28 87.81 90.13 88.05 87.10
PDS 93.83 86.77 80.84 89.49 91.82 89.66 88.73
StrategyQA SelfCheckGPT 50.86 41.74 42.72 47.37 61.67 58.73 50.52
HaloCheck 37.02 40.94 42.48 35.93 59.80 41.20 42.90
ADS 49.56 51.42 49.36 49.85 59.55 56.27 52.87
PDS 49.97 54.49 52.87 51.18 62.08 67.22 56.30
Play SelfCheckGPT 40.63 45.60 45.37 37.22 63.70 45.37 46.33
HaloCheck 37.02 40.94 42.48 35.93 59.80 41.20 42.90
ADS 44.37 48.93 49.90 46.00 72.68 50.44 52.05
PDS 46.14 48.01 51.31 47.66 72.84 51.86 52.97
Physics SelfCheckGPT 40.63 45.60 45.37 37.22 63.70 45.37 46.33
HaloCheck 37.02 40.94 42.48 35.93 59.80 41.20 42.90
ADS 39.94 40.60 48.63 79.10 53.05 52.58 52.32
PDS 46.85 48.58 61.31 78.50 56.99 60.66 58.82
FEVER SelfCheckGPT 55.08 49.36 44.86 51.63 62.28 - 52.64
HaloCheck 55.44 47.00 50.17 47.56 60.36 - 52.11
ADS 57.02 49.72 48.57 54.02 67.31 - 55.33
PDS 62.19 51.41 49.00 55.43 66.91 - 57.03
HotpotQA SelfCheckGPT 85.69 80.20 71.37 84.48 52.60 - 74.87
HaloCheck 88.06 82.18 85.20 82.51 48.20 - 77.23
ADS 89.71 79.14 83.99 90.15 74.89 - 83.58
PDS 92.61 83.21 86.21 91.47 79.41 - 86.58
2WikiMultihop SelfCheckGPT 85.69 80.20 71.37 84.48 52.60 - 74.87
HaloCheck 88.06 82.18 85.20 82.51 48.20 - 77.23
ADS 83.26 75.52 48.12 79.37 59.51 - 69.16
PDS 86.72 79.80 54.37 80.20 58.94 - 72.01
@@
Table 3: Results of PDS on HotpotQA. Metrics are F1 (%).
Method GPT3 GPT-Instruct GPT-4 Gemini Mixtral
@@
Table 4: Results of PDS on 2WikiMultihop. Metrics are F1 (%).
@@ 4 Results and Analysis
4.1 PDS Substantially Outperforms All Baselines
As shown in Table 1, Across all 45 subsets in our R2PE, PDS consistently improves ADS performance, yielding an average improvement of $5.10\%$ in terms of F1. Although ADS demonstrates greater precision, it suffers from significantly lower recall, indicating its propensity to overlook numerous examples with $L=F$. In contrast, PDS adeptly balances precision and recall, culminating in superior overall performance relative to ADS (see Table 1, 2, 17 and 18). From Table 2, PDS improves ADS in terms of AUC-PR in the majority of cases and yields comparable results (less than $1\%$ decrease) in the remaining ones. On average, PDS results in $3.01\%$ absolute improvement in AUC-PR. Notably, ADS is a fairly strong baseline compared with SelfCheckGPT and HaloCheck in our setting, and PDS can further enhance the performance of ADS. We believe that the performance of SelfCheckGPT (Manakul et al., 2023) and HaloCheck (Elaraby et al., 2023), which lags behind ADS, can be attributed to their operation in different settings and the lack of explicit verification of the majority voting proportion of the final answer, which highlights the importance of jointly considering both the reasoning process and the final answer to validate CoT, demonstrating the superiority of PDS. For discriminative tasks (StrategyQA, Play, Physics, and Play), ADS has relatively low F1 scores and AUC-PR. This may be attributed to the fixed set of answers (e.g., yes or no) in these tasks, which inherently exhibit higher consistency, making verification based solely on answer agreement challenging. In contrast, PDS can relieve this issue by evaluating reasoning processes.
@@ 4.1 PDS Substantially Outperforms All Baselines
To meticulously evaluate the efficacy of the PDS, we delve into the (HotpotQA from GPT-4) subset, partitioning it into two sets: one where answers fail to reach a consensus greater than half, and its counterpart. We then label the examples with an ADS less than the threshold $H$ as $\hat{L}=F$, and all the others as $\hat{L}=T$. As shown in Figure 3 (left), while the ADS proficiently segregates the examples with $L=T$, its discernment for the ”False” category is less acute. On the contrary, the PDS (Figure 3 middle) strikes a better balance between precision and recall. Eliminating the implicit answer verification component in Equation (1) results in diminished precision (Figure 3 right, discussed in Section 4.2).
@@ 4.2 Ablation Study
To understand the contribution of PDS’ components, we have carried out an extensive ablation study involving the following variants (See Appendix C.4):
@@
PDS w/o ans. The original PSS considers all $PPSS(R_{i},R_{j})$, where each response $R_{i}$ equals a reasoning path $r_{i}$ and an answer $a_{i}$. As the information from the answers $a_{i}$’ is already used to obtain the final answer $a$, we can discard them and only compute the entailment score among rationale pairs $PPSS(r_{i},r_{j})$, that is, $PDS_{w/o\>ans}=(\overline{ADS}+\text{mean}_{i\neq j}PPSS(r_{i},r_{j}))/2$.
@@
PDS-HaloCheck;PDS-SelfCheckNLI. HaloCheck (Elaraby et al., 2023) and SelfCheckNLI (Manakul et al., 2023) can be alternatives to PSS, denoted by PDS-halocheck and PDS-selfcheckNLI, respectively. See Appendix C.5 for discussion.
As shown in Table 4 and 4, our PDS implementation gains the best results.
@@ 4.3 PDS for Improving Downstream Performance
VE (Zhao et al., 2023) post-edits reasoning process using external knowledge (see Appendix D). VE first checks the answer agreement and then edits the reasoning chains based on a retriever if $ADS<H$. That is to say, VE selects potential incorrect samples by ADS. Our PDS can be incorporated into the VE framework, dubbed ”VE + PDS”.
Experiments are carried out on (HotpotQA from GPT-4) and (2WikiMultihop from GPT-4), utilizing ground-truth supporting contexts along with distractor paragraphs as the retrieved documents to negate the influence of the retrieval system. Our baseline is VE, and the standard prompting and CoT-SC are also included as references.
Figure 4 illustrates that integrating our method with the VE framework leads to enhanced accuracy for both datasets. Specifically, VE+PDS leads to a notable $2.9\%$ absolute increase in accuracy for HotpotQA, and $2.8\%$ for 2WikiMultihop. Given that the retrieval system in our experiments is assumed to be near-perfection, the observed improvements in performance are attributed to the more effective answer verification through the substitution of ADS with PDS. This success is likely a consequence of the PDS’s capability to supervise processes across multiple reasoning chains, leading to better selection of all potential incorrect predictions.
@@
4.4 Error Analysis
We also conduct an error analysis on HotpotQA using a GPT-4 subset of R2PE. Specifically, we find 68 cases where PDS correctly identifies the falsehood of answers that ADS does not. In these cases, all the labels are ”False” and the answers are consistent, meaning ADS could not verify the answers due to its reliance on answer consistency alone. In contrast, PDS is able to identify potential inconsistencies among different reasoning chains, even when the answers are consistent. However, we also find 53 cases where the labels are ”True” and the answers are consistent, where ADS succeeds but PDS fails. After a careful manual inspection, we classify these cases into the following categories: I. Conflicting information among chains despite consistent and correct answers. II. Excessive information that is correct in some chains, leads to inconsistency among chains. III. Imperfections of the NLI models even when there is no conflicting information. IV. Different solutions with diverging chains lead to the same correct answers. We present the number of cases in each category in Table 5. See discussion in Appendix E.
Type of Errors Type I Type II Type III Type IV
Number of Cases 14 14 18 7
Percentage (in %) 26.4 26.4 33.9 13.2
Table 5: Distribution of Error Types
@@ 5 Related Work
Extensions of CoT. CoT (Wei et al., 2022) improves LLM performance by decomposing reasoning into steps via few-shot examples. Refinements include sampling-based decoding (Wang et al., 2022), automated exemplar selection (Kojima et al., 2022; Zhang et al., 2022), and hybrid strategies (Zou et al., 2023). Task-specific adaptations like PHP (Zheng et al., 2023) for mathematical reasoning and verify-and-edit (VE) (Zhao et al., 2023) focus on improving reasoning accuracy. While much work emphasizes end-task metrics, our study explores reasoning quality-outcome linkage, verifying predictions through chain analysis. We will also extend PDS to Long-CoT LLMs (Jaech et al., 2024), which benefit from sampling-based verification (Wen et al., 2025), and propose potential extensions to test-time scaling methods (Wu et al., 2025) via chain consistency analysis.
Reasoning Chain Quality Evaluation. Rationale analysis is key for understanding AI performance and limitations(Golovneva et al., 2022; Prasad et al., 2023; He et al., 2023). Rationale evaluation methods are either reference-based, comparing against a gold standard (Clinciu et al., 2021; Saparov and He, 2022), or reference-free, using metrics like ROSCOE (Golovneva et al., 2022). While they link high-quality reasoning to better task performance, the direct impact of reasoning quality on prediction accuracy is under-researched. Our study fills this gap with a detailed quantitative analysis of this relationship. There is also research focused on evaluating not only the final predictions but also the reasoning steps in CoT (Xia et al., 2024; Huang et al., 2024; Xu et al., 2025).
Uncertainty Estimation. A range of methods measure LLM uncertainty by token probabilities (Kuhn et al., 2023; Malinin and Gales, 2020; Kadavath et al., 2022); however, they are inapplicable to R2PE, which does not furnish probabilities for responses. Moreover, they concentrate on short answer sequences, aligning more closely with the standard prompting. In contrast, our work verifies the veracity of CoT answers, which encompasses both the reasoning chains and answer sequences. Different from Kuhn et al. (2023), the entailment model in our work is used to measure the informational similarities between two responses, where the problem setting and detailed implementation are distinct.
@@ 6 Conclusion
To shed more light on the connection between the quality of the reasoning steps and the final outcome, we present R2PE, the first benchmark that quantitatively analyzes whether we can validate the answer veracity by evaluating the reasoning chains on a variety of reasoning tasks across five different domains. Six different LLMs are used in the creation of R2PE to guarantee the generalizability of our findings. We develop the PDS framework by integrating multiple rationales’ information to predict the falsehood of the final predictions, which significantly boosts the performance of answer checking on our R2PE benchmark. Furthermore, our approach can be easily combined with the verify-and-edit framework to improve end-task performance.
Acknowledgements
The work was partially supported by a grant from the Research Grants Council of the Hong Kong Special Administrative Region, China (AoE/E-601/24-N)